# Efficient absorbing boundary conditions for conservation laws: Application to the piston problem of gas dynamics

### v1

@@ 4 The Euler equations
Here $\rho$ denotes the mass density, $u$ the flow velocity, $p$ the pressure, and $\rho u$ the momentum. The energy density given by $\bar{E}=\frac{1}{2}\rho u^{2}+\rho e$, where $\rho u^{2}$ is the kinetic energy and $e$ is the specific internal energy. Equations (32a), (32b), and (32c) describe the conservation of mass, momentum, and energy respectively, and can be derived from the Navier-Stokes equations or foundational principles in continuum mechanics if viscosity and heat flux are neglected [24].
We deal with an ideal gas with constant specific heats, hence we can define the constant heat capacity ratio $\gamma=c_{p}/c_{V}$, and our equation of state becomes [41]
@@
Throughout this work, we assume the density $\rho$ to be strictly positive for all times. Therefore vacuum states, which require special analytical and numerical treatment [38, Sec. 4.6], are not considered.
It is often convenient to write Equations (32) in Lagrangian form. Given a real function $\tilde{\xi}(\cdot)$, we will keep track of a set of conserved quantities at each particle defining the trajectory map [35]
@@
where $E(x,t)=\bar{E}/\rho$ [37]. Furthermore, using the equation of state (33), we have the closure relation
@@
where $\mathbf{q}=\left(q_{1},q_{2},q_{3}\right)^{T}=\left(V,u,E\right)^{T}$ and $\mathbf{f}(\mathbf{q})=\left(q_{1},p(\mathbf{q}),q_{2}p(\mathbf{q})\right)^{T}=\left(-u,p,up\right)^{T}$ denote the vector of conserved quantities and the nonlinear flux function respectively. For smooth solutions, System (39) can be written in quasilinear form
@@
has symmetric eigenvalues $\lambda_{1}(\mathbf{q})=-\sqrt{\gamma p/V}$, $\lambda_{2}=0$, $\lambda_{3}(\mathbf{q})=\sqrt{\gamma p/V}$, with corresponding eigenvectors $\mathbf{r}_{1,3}(\mathbf{q})=\left(-\frac{\sqrt{V}}{\sqrt{p}},\mp\sqrt{\gamma},\sqrt{pV}\mp u\sqrt{\gamma}\right)^{T},\,\mathbf{r}_{2}(\mathbf{q})=\left(\gamma-1,0,p\right)^{T}.$
@@
has eigenvalues $\lambda_{1}=-\gamma$, $\lambda_{2}=0$, $\lambda_{3}=\gamma$, and eigenvectors $\mathbf{r}_{1}=\left(-\frac{1}{\gamma},-1,1\right)^{T},$ $\mathbf{r}_{2}=\left(-\frac{1-\gamma}{\gamma},0,1\right)^{T},$ $\mathbf{r}_{3}=\left(-\frac{1}{\gamma},1,1\right)^{T}.$
@@ 4.1 The piston problem
Let us formulate the piston problem of gas dynamics [15], which will be the main application for the numerical treatment of boundary conditions studied in this work. The equations (32) are solved in the domain $\Omega_{t}=[\tilde{\xi}(t),+\infty[$ with a constant steady state $\overline{\mathbf{q}}=(1,0,\gamma/(\gamma-1))$ as initial condition. The left physical boundary $\tilde{\xi}(t)$ corresponds to the position of an oscillatory plate given by
[…]
@@
In the recent work [35, Ch. 3], Phan Thi My studies the treatment of outflow boundary conditions for this problem. There, the author replaces the Euler equations by a Burgers-type equation in some additional ghost cells. The simple wave behavior of the solution in this region greatly diminishes the unphysical reflections with just a few ghost cells. While using a fixed small number of ghost cells is desirable from an efficiency standpoint, it does not allow for better absorption by using a larger sponge layer.
@@ 5 Numerical results
We analyze the performance of the ABCs described in the previous sections for the nonlinear equations (39) and their linearized version (42). The setup described at the beginning of Section 2 is considered, where the inflow boundary condition $\mathbf{q}_{BC}$ is given by the piston’s motion.
5.1 Discretization details and setup
For the numerical results presented in this work, we use a high-order wave propagation formulation of Godunov-type methods, available in the Clawpack software package [29] and accessible through the Pyclaw framework [21]. For a detailed description of these methods, the reader is referred to the works [22, 20].
Denoting by $Q_{i}$ the vector of conserved quantities $\mathbf{q}$ averaged over the volume $[x_{i-1/2},x_{i+1/2}]$, our semidiscretization of Systems (39) and (42) reads
@@ 5.1 Discretization details and setup
The fluctuations $\mathcal{A}^{\pm}\Delta q_{i-\frac{1}{2}}$ represent the net contribution of the left and right-going waves emanating from the solution of a Riemann problem at the interface $x_{i-1/2}$. The initial data for each Riemann problem is given by a piecewise polynomial reconstruction of the solution at both sides of the interface. To avoid spurious oscillations around strong shocks without using an extremely refined grid, a piecewise linear reconstruction with the TVD minmod limiter is used. An advantage of the semi-discretization (46) is that it can be applied to systems in nonconservative form, like (3), in a natural way [2, 22]. Since the semidiscretization and numerical inflow boundary condition (see Equations (48) and (49) below) formally attain second order of accuracy, we use Heun’s two-stage second-order accurate Runge-Kutta scheme to integrate System (46) in time.
To compute an approximate Riemann solution at an interface with left and right constant initial data $q_{L},\,q_{R}$, we use an HLL Riemann solver [11]. The approximate speeds are defined to bound the wave speeds from the exact Riemann solution from below and above,
@@
Since the eigenvalues of our Jacobian matrix are symmetric, we have $s^{1}=-s^{2}$. Therefore, in this case, the HLL solver coincides with the Local Lax-Friedrichs (or Rusanov) solver.
The computation of fluctuations and the reconstruction process require the use of ghost cells at the left and right boundaries of the domain, denoted by $\mathcal{C}_{-1}$ and $\mathcal{C}_{N_{x}}$ respectively. We assume the piston (of infinitesimal width) to be located at the interface between the cells $\mathcal{C}_{-1}$ and $\mathcal{C}_{0}$ ($x=0$). Using the flow conditions at the piston, formulated in [7], and their discretization described in [35], we get the second-order accurate approximations
@@
$V^{[n]}_{-1}=\frac{1}{\gamma}(p^{[n]}_{0}-p^{[n]}_{-1}+\gamma V^{[n]}_{0}),\quad E^{[n]}_{-1}=\frac{1}{\gamma-1}(p^{[n]}_{-1}+\gamma V^{[n]}_{-1}),$ (49b)
where we have the closure relation $p=\gamma V+(\gamma-1)E$ instead of (38). At the right boundary we set $Q_{N_{x}}=\overline{\mathbf{q}}$, where $\overline{\mathbf{q}}$ is the far-field state $\overline{\mathbf{q}}=(1,0,\gamma/(\gamma-1))$ for the nonlinear equations (39), and $\overline{\mathbf{q}}=\bf 0$ for the linear equations (42). Here, $\mathcal{C}_{N_{x}}$ is the ghost cell centered at $x_{\infty}$. For both systems, the initial condition will be their corresponding far-field state $\overline{\mathbf{q}}$.
For all the numerical tests below, $x_{s}=20\pi$, and $x_{\infty}$ takes different values to evaluate the dependence of the ABCs on the sponge layer’s length $\omega=|x_{\infty}-x_{s}|$. Since the piston’s frequency is $(2\pi)^{-1}$, as pointed out in Equation (44), the characteristic wavelength of the solution will be $L=2\pi$. Different spatial refinement levels are considered, such that $N$ cells are contained in a wavelength $L$, for $N\in\{10,50,250\}$. A Courant number $C=0.8$ is used to compute $\Delta t$ at each time step under the standard CFL condition $\Delta t={C\Delta x}/{\max_{i}\max_{p}{|\lambda_{p}(Q_{i})|}}.$ The amplitude of the piston’s oscillations is set to $M=0.4$, which in the nonlinear case will lead to a solution dominated by strong shocks, as studied by Inoue and Yano [15].
@@ 5.2 Parameter configuration
  • a slowing-down ($s$) and a damping ($d$) function for the far-field operators-based techniques (SDO, S-SDO, NDO),
  • a scalar weight function ($\Gamma$) for the relaxation methods (RM, RM-M).
These functions are constructed using $\Gamma_{A}$ (9) and $\Gamma_{B}$ (with $b=1/2$) (11), by taking
$\Gamma(x)=\Gamma_{(\cdot)}(x),\quad s_{(\cdot)}(x)=\Gamma_{(\cdot)},\quad d_{(\cdot)}(x)=\sigma\left(1-\Gamma_{(\cdot)}(x)\right),$ (52)
where $\sigma$ is the maximum damping rate attained in the sponge layer. The shapes of these functions are illustrated in Figure 3.
@@
First, we evaluate different combinations of $s(\cdot)$ and $d(\cdot)$ for SDO with $\sigma=30$, $N=250$, and considering different sponge layer lengths $\omega$. The relative errors obtained with each combination are displayed in Table 1. For the linear equations (42), while the choice of $s$ and $d$ does not seem to be critical, the best performance is attained with $s_{A}$ and $d_{B}$. With this configuration, the slowing-down effect is greater than the damping at the early stage of the absorbing process. Thus, the length of the outgoing waves is reduced before they are damped and, consequently, the drawbacks associated with large wavelengths, addressed in Section 2, are alleviated.
{tblr}
cells = c, column3 = Mercury, column8 = Mercury, cell11 = r=2, cell12 = c=4, cell16 = c=4, vlines, hline1,3,7 = -, hline2 = 2-9, rowsep = 1pt, $\frac{\omega}{L}$ Linear equations (42) Nonlinear equations (39)
@@
In contrast with the linear scenario, the choice of slowing-down and damping functions drastically affects the behavior of the SDO ABC for the nonlinear equations (39) (see Table 1). For this system, $s_{A}$ and $d_{B}$ yield the least error overall. This can be attributed to the slowing-down operator’s tendency to generate entropy inside the sponge layer, illustrated in Appendix A.3, which needs to be dissipated by a rapidly increasing damping rate. The effect of using different maximum damping rates $\sigma$ with SDO is depicted in Figure 3. Since no significant improvement is obtained for larger values, $\sigma=30$ is used for the SDO ABC in the following. Proceeding analogously, we select $d_{B}$ and $\sigma=20$ for the NDO ABC, as suggested by the results displayed in Table 5 and Figure 5.
@@
{tblr}
cells = c, column3 = Mercury, cell11 = r=2, cell12 = c=2, vlines, hline1,3,7 = -, hline2 = 2-3, rowsep = 1pt, $\frac{\omega}{L}$ Nonlinear equations (39)
@@
$1$ 3.21e-03 2.03e-03
Figure 4: Relative error (50) for the nonlinear equations (39) with NDO ABC 3.4 and different damping functions $(\sigma=20,\,N=250)$
@@
{tblr}
cells = c, column3 = Mercury, column5 = Mercury, column7 = Mercury, column9 = Mercury, cell11 = r=3, cell12 = c=4, cell16 = c=4, cell22 = c=2, cell24 = c=2, cell26 = c=2, cell28 = c=2, vlines, hline1,4,8 = -, hline2-3 = 2-9, rowsep = 1pt, $\frac{\omega}{L}$ Linear equations (42) Nonlinear equations (39)
@@ 5.3 Performance comparison
Let us compare the behavior of the different relaxation methods discussed in Sections 2.2, 3.2, and 3.3. We refer to Algorithms 1 and 2 as RM and RM-M respectively, and to their use described in Algorithm 3 as RM 2 and RM-M 2. For the latter, the weight function $\Gamma_{B}^{2}$ is used, according to the observations from Section 3.3. Moreover, we consider the application of these techniques after each stage of the Runge-Kutta time integration, as proposed by Wasistho et al. [39], and refer to them as RM RK and RM-M RK. The results are displayed in Figures 7 and 7 for the linear (42) and nonlinear (39) equations respectively, along with the errors corresponding to the numerical discretization (51) and the use of constant extrapolation at the right boundary.
For the linear equations, there is no perceivable difference between RM and RM-M, and their variants. This can be attributed to the fact that due to the piston motion, all the waves that enter the sponge layer are right-going, and no interaction between them takes place in the linear setting. Therefore, there is no difference betwen absorbing waves selectively or not. Conversely, for the nonlinear equations, there is a perceivable improvement with the RM-M with respect to the RM when larger portions of a characteristic wavelength $L$ are contained in the sponge layer. The techniques RM and RM-M attain exactly the same error as RM 2 and RM-M 2 respectively for both systems. This validates the equivalence between Algorithms 2 and 3, established in Section 3.3 for the linear equations, and suggests that it extends to the nonlinear case. Additionally, as the grid is refined, we see that the error obtained with RM-M-RK decays in a smoother than the other relaxation methods as the sponge layer length $\omega$ increases.
@@
Now we assess the perfomance of all the ABCs addressed in Sections 2 and 3; the results are displayed in Figures 9 and 9. For the linear equations, the behavior of the different ABCs is similar for all the grid sizes, with the exception of S-SDO, which attains a larger error that diminishes as the grid is refined.
@@
For the nonlinear equations, the methods based on far-field operators (SDO, S-SDO, NDO) fail to provide a satisfactory error and the numerical solution diverges for the coarsest grids. This suggests that the discretization is not stable with the time step size prescribed by the CFL condition with the chosen Courant number. Indeed, we can see that the $L^{2}$ norm of $\mathbf{D}(x)$ in Equation (3) is $\max_{i}|d_{i}s_{i}\lambda_{i}|$, which is larger than time step size restriction for a coarse grid when the maximum damping parameter $\sigma$ is large. As seen in Figure 3, a large maximum damping rate might be necessary to attain a prescribed error for a fixed sponge layer length. Since it is convenient to use an explicit scheme for the hyperbolic part of the problem, a way to deal with this artificial stiffness, not explored in this work, is to use an IMEX scheme [33] to evolve the semidiscretization. On the other hand, the RM and its extension RM-M, introduced herein, show to be the most robust ABCs for both the linear and nonlinear equations.
@@
Once the grid (and thus the time step size) is refined, and the stability issues are overcome, we can see that the SDO ABC attains its minimum error for a sponge layer length of around $0.08L$. It is followed by RM-M and NDO which attain their minimum error for sponge layer lengths of $0.6L$ and $5L$ respectively. While the RM attains a low error for a sponge layer length of $0.04L$, it still does not reach the absborption of the other methods, even for a sponge layer length of $10L$. In our particular setting, we know that the first two shocks of the wave train are periodically merged in the far-field [15], and such interactions might cause information to propagate back into the computational domain [12]. Consequently, as the sponge layer size increases, methods that selectively damp a characteristic field, like SDO, NDO, and RM-M, allow this physical information to recirculate, while scalar techniques, like RM and S-SDO, fail to capture this behavior (see Remark 1 below).
@@
Finally an execution time comparison between the different ABCs for the nonlinear equations (39) is displayed in Table 3. We consider the best wall clock time obtained for each method after 7 runs on a MacBook Pro 2019 with a 2.3 GHz 8-Core Intel Core i9 processor. As expected, the scalar methods, RM and S-SDO, are the fastest ones, followed by the directional methods RM-M SDO. The NDO technique is, by a large margin, the least efficient due to the computationally expensive treatment of its nonlocal integral term (see Appendix A.2).
Remark 1.
To further illustrate the effect of damping waves corresponding to a particular characteristic field, we consider an artificial setup where an incoming pulse is created inside the sponge layer as shown in Figure 10. The initial pulse is immediately damped by the scalar methods (RM and S-SDO), while just the right-going components of it are absorbed by the directional methods (RM-M and SDO).
Figure 10: Above: Effect of directional (SDO) and scalar (SSDO) damping operators applied to an incoming physical wave. The sponge layer is colored light grey. Below: Effect of directional relaxation method with matrix-valued weight function (RM-M) and scalar relaxation method (RM) applied to an incoming physical wave.
@@
10 0.356278 0.413172 0.311402 0.338358 0.487306
50 1.37656 1.89572 1.49388 1.32294 4.84846
250 14.7071 31.6987 10.8559 32.3652 107.529
Table 3: Execution time comparison of ABCs with different refinement levels $N$ for the nonlinear equations with a sponge layer of length $\omega=10L$

### v2

@@ 4 The Euler equations
Here $\rho$ denotes the mass density, $u$ the flow velocity, $p$ the pressure, and $\rho u$ the momentum. The energy density given by $\bar{E}=\frac{1}{2}\rho u^{2}+\rho e$, where $\rho u^{2}$ is the kinetic energy and $e$ is the specific internal energy. Equations (38a), (38b), and (38c) describe the conservation of mass, momentum, and energy respectively, and can be derived from the Navier-Stokes equations or foundational principles in continuum mechanics if viscosity and heat flux are neglected [landau_ideal_1987].
We deal with an ideal gas with constant specific heats, hence we can define the constant heat capacity ratio $\gamma=c_{p}/c_{V}$, and our equation of state becomes [whitham_linear_1974]
@@
Throughout this work, we assume the density $\rho$ to be strictly positive for all times. Therefore vacuum states, which require special analytical and numerical treatment [toro_riemann_2009, Sec. 4.6], are not considered.
@@
where $E(x,t)=\bar{E}/\rho$ [smoller_shock_1994]. Furthermore, using the equation of state (39), we have the closure relation
@@
where $\mathbf{q}=\left(q_{1},q_{2},q_{3}\right)^{T}=\left(V,u,E\right)^{T}$ and $\mathbf{f}(\mathbf{q})=\left(q_{1},p(\mathbf{q}),q_{2}p(\mathbf{q})\right)^{T}=\left(-u,p,up\right)^{T}$
denote the vector of conserved quantities and the nonlinear flux function respectively. For smooth solutions, System (45) can be written in quasilinear form
@@
has symmetric eigenvalues $\lambda_{1}(\mathbf{q})=-\sqrt{\gamma p/V}$, $\lambda_{2}=0$, $\lambda_{3}(\mathbf{q})=\sqrt{\gamma p/V}$, with corresponding eigenvectors $\bm{r}_{1,3}(\bm{q})=\left(-\frac{\sqrt{V}}{\sqrt{p}},\mp\sqrt{\gamma},\sqrt{pV}\mp u\sqrt{\gamma}\right)^{T},\,\bm{r}_{2}(\bm{q})=\left(\gamma-1,0,p\right)^{T}.$
@@
has eigenvalues $\lambda_{1}=-\gamma$, $\lambda_{2}=0$, $\lambda_{3}=\gamma$, and eigenvectors $\bm{r}_{1}=\left(-\frac{1}{\gamma},-1,1\right)^{T},$ $\bm{r}_{2}=\left(-\frac{1-\gamma}{\gamma},0,1\right)^{T},$ $\bm{r}_{3}=\left(-\frac{1}{\gamma},1,1\right)^{T}.$
@@ 4.1 The piston problem
Let us formulate the piston problem of gas dynamics [inoue_propagation_1993], which will be the main application for the numerical treatment of boundary conditions studied in this work.
The equations (38) are solved in the domain $\Omega_{t}=[\tilde{\xi}(t),+\infty[$ with a constant steady state $\overline{\mathbf{q}}=(1,0,\gamma/(\gamma-1))$ as initial conditions. The left physical boundary $\tilde{\xi}(t)$ corresponds to the position of an oscillatory plate given by
$\tilde{\xi}(t)=M\left(\cos(t)-1\right)\ \mathbbm{1}_{[\mathbb{R}^{+}]}(t),$ (50)
@@
In the recent work [phan_thi_my_finite_2021, Ch. 3], Phan Thi My investigates outflow boundary conditions for this problem by replacing the Euler equations with a Burgers-type model in the sponge layer. The simple wave behavior reduces spurious reflections even when the sponge layer contains only a small number of cells (fewer than ten). This approach is attractive because it achieves efficient absorption with minimal computational overhead. However, its performance does not improve when the sponge layer is thickened, which limits its flexibility.
@@ 5 Numerical results
We analyze the performance of the ABCs introduced in the previous sections for the nonlinear equations (45) and their linearized version (48). For this, we consider the setup described at the beginning of Section 2. Throughout this section, ”performance” refers to the effectiveness of an ABC in minimizing wave reflections, quantified by the reflection error (56).
5.1 Discretization details
For the numerical results presented in this work, we use a high-order wave propagation formulation of Godunov-type methods, available in the Clawpack software package [mandli_clawpack_2016] and accessible through the Pyclaw framework [ketcheson_pyclaw_2012]. For a detailed description of these methods, the reader is referred to the works [ketcheson_high-order_2013, ketcheson_wenoclaw_2008].
Denoting by $Q_{i}$ the vector of conserved quantities $\mathbf{q}$ averaged over the volume $x_{i-1/2},x_{i+1/2}]$, our semidiscretization of Systems (45) and (48) reads
@@ 5.1 Discretization details
The fluctuations $\mathcal{A}^{\pm}\Delta q_{i-\frac{1}{2}}$ represent the net contribution of the left and right-going waves emanating from the solution of a Riemann problem at the interface $x_{i-1/2}$. The initial data for each Riemann problem is given by a piecewise polynomial reconstruction of the solution at both sides of the interface. To avoid spurious oscillations around strong shocks without using an extremely refined grid, a piecewise linear reconstruction with the TVD minmod limiter is used. An advantage of the semi-discretization (52) is that it can be applied to systems in nonconservative form, like (3), in a natural way [bale_wave_2003, ketcheson_high-order_2013]. Since the semidiscretization and numerical inflow boundary condition (see Equations (54) and (55) below) formally attain second order of accuracy, we use Heun’s two-stage second-order accurate Runge-Kutta scheme to integrate System (52) in time. A Courant number $C=0.8$ is used to compute $\Delta t$ at each time step under the standard CFL condition $\Delta t={C\Delta x}/{\max_{i,p}{|\lambda_{p}(Q_{i})|}}.$
To compute an approximate Riemann solution at an interface with left and right constant initial data $q_{L},\,q_{R}$, we use an HLL Riemann solver [harten_upstream_1983]. The approximate speeds are defined to bound the wave speeds from the exact Riemann solution from below and above,
@@
Since the eigenvalues of our Jacobian matrix are symmetric, we have $s^{1}=-s^{2}$. Therefore, in this case, the HLL solver coincides with the Local Lax-Friedrichs (or Rusanov) solver. Since the solutions we are focused on do not develop vacuum states, the use of the HLL solver helps to avoid their spurious appearance.
We now turn our attention to the piston problem described in Section 4.1, which will be used as a benchmark to evaluate the performance of the different ABCs. Let us denote the $i$-th cell of the discretized domain by $\mathcal{C}_{i}=[x_{i-1/2},x_{i+1/2}]$, and the ghost cells at the left and right boundaries by $\mathcal{C}_{-1}$ and $\mathcal{C}_{N_{x}}$ respectively. We assume the piston (of infinitesimal width) to be located at the interface between the cells $\mathcal{C}_{-1}$ and $\mathcal{C}_{0}$ ($x=0$). Using the flow conditions at the piston, formulated in [fazio_central_2010], and their discretization described in [phan_thi_my_finite_2021], we get the second-order accurate approximations
@@
[…]
where we have the closure relation $p=\gamma V+(\gamma-1)E$ instead of (44). At the right boundary we set $Q_{N_{x}}=\overline{\mathbf{q}}$, where $\overline{\mathbf{q}}$ is the far-field state $\overline{\mathbf{q}}=(1,0,\gamma/(\gamma-1))$ for the nonlinear equations (45), and $\overline{\mathbf{q}}=\bf 0$ for the linear equations (48). Here, $\mathcal{C}_{N_{x}}$ is the ghost cell with left boundary $x_{\infty}$. For both systems, the initial condition will be their corresponding far-field state $\overline{\mathbf{q}}$.
For all the numerical tests below, $x_{s}=20\pi$, and $x_{\infty}$ takes different values to evaluate the dependence of the ABCs on the sponge layer’s length $\omega=|x_{\infty}-x_{s}|$. Since the piston’s frequency is $(2\pi)^{-1}$, as pointed out in Equation (50), the characteristic wavelength of the solution will be $L=2\pi$. Different spatial refinement levels are considered, such that $N$ cells are contained in a wavelength $L$, for $N\in\{10,50,250\}$. The amplitude of the piston’s oscillations is set to $M=0.4$, which in the nonlinear case will lead to a solution dominated by strong shocks, as studied by Inoue and Yano [inoue_propagation_1993].
@@ 5.2 Parameter configuration
  • the slowing-down and damping functions $s$ and $d$, and the damping rate $\sigma$ (5) for the construction of the slowing-down and damping far-field operators (SDO, S-SDO, NDO),
  • a scalar weight function $\Gamma$ for the relaxation methods (RM, RM-M).
The functions $s$, $d$, and $\Gamma$ will be constructed using either $\Gamma_{A}$ (15) or $\Gamma_{B}$ (with $b=1/2$) (17). The slowing-down and damping functions constructed with $\Gamma_{A}$ ($\Gamma_{B}$) will be denoted by $s_{A}$ ($s_{B}$) and $d_{A}$ ($d_{B}$) respectively. The shapes of these functions are illustrated in Figure 4.
@@
First, we evaluate different combinations of $s$ and $d$ for SDO with $\sigma=30$, $N=250$, and consider different sponge layer lengths $\omega$. The relative errors obtained with each combination are displayed in Table 1. For the linear equations (48), while the choice of $s$ and $d$ does not seem to be critical, the best performance is attained with $s_{A}$ and $d_{B}$. With this configuration, the slowing-down effect is greater than the damping at the early stage of the absorbing process. Thus, the length of the outgoing waves is reduced before they are damped and, consequently, the drawbacks associated with large wavelengths, addressed in Section 2, are alleviated.
$\frac{\omega}{L}$ Linear equations (48) Nonlinear equations (45)
@@
In contrast with the linear scenario, the choice of slowing-down and damping functions drastically affects the behavior of the SDO ABC for the nonlinear equations (45) (see Table 1). For this system, $s_{B}$ and $d_{A}$ yield the least error overall. This can be attributed to the slowing-down operator’s tendency to generate entropy inside the sponge layer, illustrated in Appendix D, which needs to be dissipated by a rapidly increasing damping rate. The effect of using different maximum damping rates $\sigma$ with SDO is depicted in Figure 4. Since no significant improvement is obtained for larger values, $\sigma=30$ is used for the SDO ABC in the following. Proceeding analogously, we select $d_{B}$ and $\sigma=20$ for the NDO ABC, as suggested by the results displayed in Table 2 and Figure 5.
@@
Although some of these parameters could be optimized through reflection-coefficient-based strategies [peric2018analytical], in this work we tune them based on the error (56) only to ensure a fair and consistent performance comparison among the different ABCs.
$\frac{\omega}{L}$ Nonlinear equations (45)
@@
11 3.21e-03 2.03e-03
Table 2: Relative error (56) for the nonlinear equations (45) with NDO ABC 3.4 and different damping functions $(\sigma=20,\,N=250)$
@@
$\frac{\omega}{L}$ Linear equations (48) Nonlinear equations (45)
@@ 5.3 Performance comparison
Let us compare the behavior of the different relaxation methods discussed in Sections 2.2, 3.2, and 3.3. We refer to Algorithms 1 and 2 as RM and RM-M respectively, and to their use described in Algorithm 3 as RM 2 and RM-M 2. For the latter, the weight function $\Gamma_{B}^{2}$ is used, according to the observations from Section 3.3. Moreover, we consider the application of these techniques after each stage of the Runge-Kutta time integration, as proposed by Wasistho et al. [wasistho_simulation_1997], and refer to them as RM RK and RM-M RK. The results are displayed in Figures 7 and 7 for the linear (48) and nonlinear (45) equations respectively.
To isolate the effect of the ABCs, the error associated with the numerical discretization (57) is displayed in black dot-dashed lines with triangle markers in all the figures of this section. Furthermore, for the sake of comparison, the error associated with the use of constant extrapolation at the last ghost cell without a sponge layer (a simple and common boundary treatment) is shown in grey dotted lines with star markers.
For the linear equations, there is no perceivable difference between RM and RM-M, and their variants. This can be attributed to the fact that due to the piston motion, all the waves that enter the sponge layer are right-going, and no interaction between them takes place in the linear setting. Therefore, there is no difference between absorbing waves selectively or not. Conversely, for the nonlinear equations, the only noticeable difference is between the scalar (RM) and directional variants (RM-M) of the relaxation method when larger portions of a characteristic wavelength $L$ are contained in the sponge layer.
@@
Now we assess the perfomance of all the ABCs addressed in Sections 2 and 3; the results are displayed in Figures 9 and 9. The methods based on far-field operators (SDO, S-SDO, NDO) fail to provide a satisfactory error and the numerical solution diverges for the coarsest grids. This suggests that the discretization is not stable with the time step size prescribed by the CFL condition with the chosen Courant number. Indeed, we can see that the $L^{2}$ norm of $\mathbf{D}(x)$ in Equation (3) is $\max_{i}|d_{i}s_{i}\lambda_{i}|$, which is larger than time step size restriction for a coarse grid when the maximum damping parameter $\sigma$ is large. As seen in Figure 4, a large maximum damping rate might be necessary to attain a prescribed error for a fixed sponge layer length. Since it is convenient to use an explicit scheme for the hyperbolic part of the problem, a way to deal with this artificial stiffness, not explored in this work, is to use an ImEx scheme [pareschi_implicit-explicit_2005] to evolve the semidiscretization. On the other hand, the RM and its extension RM-M, introduced herein, show to be the most robust ABCs for both the linear and nonlinear equations.
@@
Once the grid (and thus the time step size) is refined, and the stability issues are overcome, we can see that the SDO ABC attains its minimum error for a sponge layer length of around $0.08L$. It is followed by RM-M and NDO which attain their minimum error for sponge layer lengths of $0.6L$ and $5L$ respectively. While the scalar methods S-SDO and RM attain a low error for a sponge layer length of $0.04L$, they still do not reach the absorption of the other methods, even for a sponge layer length of $10L$. Indeed, despite the fact that all the ABCs eventually attain an error that is approximately $1.5$ orders of magnitude smaller than the numerical discretization error, the directional methods (SDO, RM-M) further improve upon the scalar ones (S-SDO, RM) by about half an order of magnitude for the nonlinear equations. This can be explained by the fact that, in the nonlinear setting, the first two shocks of the wave train are periodically merged in the far-field [inoue_propagation_1993], and such interactions might cause information to propagate back into the computational domain [hedstrom_nonreflecting_1979]. Consequently, as the sponge layer size increases, methods that selectively damp a characteristic field, like SDO, NDO, and RM-M, allow this physical information to recirculate, while scalar techniques, like RM and S-SDO, fail to capture this behavior. Even though a sponge layer length much larger than the characteristic wavelength $L$ is not generally used in practice, it is useful in Figures 9 and 9 to illustrate that a larger sponge layer does not compensate for the lack of selectivity of scalar methods.
Although unphysical values (e.g. negative pressure) are not observed with any method, it is worth mentioning that, being a convex combination of admissible states, the RM is the only method that is a priori guaranteed to preserve physical admissibility of the solution (as long as the underlying numerical scheme does so). For close-to-vacuum states, further care might be required for the other methods to maintain physical values.
@@
Finally an execution time comparison between the different ABCs for the nonlinear equations (45) is displayed in Table 4. We consider the best wall clock time obtained for each method after 7 runs on a MacBook Pro 2023 with a 10-Core CPU Apple M2 Pro chip. As expected, the scalar methods, RM and S-SDO, are the fastest ones, closely followed by the directional methods RM-M and SDO. The NDO technique is, by a large margin, the most computationally expensive due to the treatment of its nonlocal integral term (see Appendix C).
@@
10 0.12 0.14 0.12 0.14 0.15
50 0.56 0.79 0.57 0.66 1.78
250 6.91 11.72 8.12 11.40 40.96
Table 4: Execution time (s) comparison of ABCs with different refinement levels NN for the nonlinear equations with a sponge layer of length $\omega=10L$